Circular No. 8398

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2004 R1 R. H. McNaught reports his discovery of a comet on CCD images taken

by himself with the 0.5-m Uppsala Schmidt telescope at Siding Spring on Sept. 2.4 UT, when the object was diffuse with a hint of a tail to the southeast. Images taken by McNaught with the Siding Spring 1.0-m f/8 reflector on Sept. 4.4 show the comet to be diffuse with a 10′′ coma and a tail 15′′ long in p.a. 110^{o}. Available astrometry by McNaught:

2004 UT  α_{2000}  δ_{2000}  Mag.
Sept. 2.42449  15^{h}06^{m}39.^{s}13  −16^{o}52′53′′.6  17.6
2.43392  15 06 42.02  −16 52 58.7  17.5
2.44340  15 06 44.90  −16 53 04.8  17.4
2.45293  15 06 47.63  −16 53 09.9  18.3
4.36913  15 16 28.16  −17 11 24.4  19.1
4.37098  15 16 28.75  −17 11 24.9  19.1
4.37284  15 16 29.20  −17 11 25.8  19.4
4.42519  15 16 45.15  −17 11 54.0  19.2

COMETS C/2002 K11 AND C/2004 P2–P5 (SOHO) Continuation to IAUC 8394 (all Kreutz sungrazers; HS = H. Su):

Comet  2002 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2002 K11 May 26.671 4^{h}14.^{m}7 +19^{o}39′ C2 DE 2004-Q11
Comet  2004 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2004 P2 Aug. 4.096 8^{h}41.^{m}6 +15^{o}56′ C3 DE 2004-R02
C/2004 P3  5.638 8 46.6 +15 36 C3 JS 2004-R02
C/2004 P4  6.346 8 50.3 +15 21 C3 XL 2004-R02
C/2004 P5  8.571 8 55.7 +13 21 C3 HS 2004-R02

COMET C/2004 Q2 (MACHHOLZ) Visual total-magnitude and coma-diameter estimates in moonlight:

Aug. 28.16 UT, 10.9, 2′.3 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector); 30.88, 11.0, 1′.8 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 31.71, 10.5:, ∼ 2′.0 (J. Drummond, Gisborne, New Zealand, 0.41-m reflector).

2004 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green